OpenStudy (anonymous):

the combined land area of the countries a and b is 159,973 square kilometers.country a is larger by 373 square kilometers. determine the land area of each country.

OpenStudy (anonymous):

Country A = x Country B = x - 373 Combination = 159,973 Can you do the rest?

OpenStudy (anonymous):

x + (x - 373) = 159,973

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

x + (x - 373) = 159973 2x - 373 = 159973 2 x = 160346 Got it now?

OpenStudy (anonymous):

x=80173

OpenStudy (anonymous):

\[x = 80,173\] Now you can find country B squared kilometers! \[x - 373 \] \[80,173 - 373\] \[= 78,900 \]

OpenStudy (anonymous):

ok 78,900

OpenStudy (anonymous):

79,900

OpenStudy (anonymous):

i got big fat red x

OpenStudy (anonymous):

Did I write it wrong?

OpenStudy (anonymous):

79,800sq. kilo

OpenStudy (anonymous):

80,173 and 79,900

OpenStudy (anonymous):

I switched the two numbers. My apologies!

OpenStudy (anonymous):

i got a new it same but new number

OpenStudy (anonymous):

the combined land area of the countries a and b is 193,973 square kilometers.country a is larger by 873 square kilometers. determine the land area of each country.

OpenStudy (anonymous):

Well, its the same method. Sorry that I wrote it wrong!

OpenStudy (anonymous):

it ok i got 3 try

OpenStudy (anonymous):

193,973 = x + (x - 873) 193,973 = 2x - 873 194,846 = 2x x = 97423

OpenStudy (anonymous):

Check my work to see if I got it! :)

OpenStudy (anonymous):

x - 873 = ? = 97423 - 873 = 96550

OpenStudy (anonymous):

im lost now

OpenStudy (anonymous):

ilfy214 lost

OpenStudy (anonymous):

why? One is 97423 and the other is 96550...

OpenStudy (anonymous):

im not lost

OpenStudy (anonymous):

I'm lost?

OpenStudy (anonymous):

im dumb u did good

OpenStudy (anonymous):

oh okay! :) Well, I'm gonna go now.... Lolz

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem!